# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ Abstract
Foundation models have shown remarkable capabilities across diverse multi-modal tasks, but their centralized training raises privacy concerns and induces high transmission costs. In contrast, federated learning (FL) offers a distributed alternative without the need to share data. Recently, for the growing demand for personalizing AI models for different user purposes, personalized federated learning (PFL) has emerged. PFL allows each client to leverage the knowledge of other clients for further adaptation to individual user preferences, again without the need to share data. Despite its potential, most PFL studies remain confined to simulated environments, overlooking the data and model heterogeneity that arise in real-world scenarios. In contrast, we first consider large data heterogeneity, evaluating on a new benchmark for multi-modal PFL, spanning 40 distinct tasks with realistic data distribution shifts. We then consider model heterogeneity in that we do not assume that all clients share similar model architectures. To address data heterogeneity, we propose a task-similarity-aware model aggregation method that provides customized global models to each client. For model heterogeneity, we propose a dimension-invariant module that enables knowledge sharing across heterogeneous models. Empirical validations demonstrate that the proposed approach outperforms the state-of-the-art, excelling in both personalization and generalization capabilities.
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) with billions of parameters, such as GPT-4o (hurst2024gpt, ), Qwen2.5-VL (yang2024qwen2, ) and Gemini (team2024gemini, ), have recently demonstrated remarkable performance. This is driven by training on large-scale data collected from various sources using high-performance computing resources, which allows for strong generalizability across diverse multi-modal tasks (chen2024feddat, ). However, such centralized training raises significant challenges, including concerns about data privacy and high transmission costs (alvi2022federated, ; MIR-2022-03-072, ). Moreover, as the demand for personalization grows, further fine-tuning is required to adapt models to individual user preferences (lau2024personalized, ; zhang2024guided, ).
To address both privacy and personalization challenges, personalized federated learning (PFL) (smith2017federated, ) has emerged as a decentralized alternative. The PFL shares locally trained models instead of raw data, while leveraging the knowledge of other clients to further adapt each local model to individual user preferences. Recent PFL methods show that the sharing of knowledge among client models, trained in a distributed manner, significantly improves the performance of personalized MLLMs (scott2024pefll, ; xie2024perada, ).
Figure 1: Overview of Heterogeneous PFL. $L_{i}$ refers to the local model for the $i_{th}$ client. For each client, new data are encountered continuously in a streaming manner, with each client focusing on a different task (i.e., data heterogeneity). In addition to data heterogeneity, model architectures may differ across clients (i.e., model heterogeneity) due to differences in hardware constraints.
@@
Despite its potential, most PFL studies remain confined to simulated environments, as they often overlook the client heterogeneity present in real-world scenarios (bujotzek2025real, ; zhang2024fedcvd, ). Specifically, in real-world scenarios, clients typically have distinct models tailored to their environments (i.e., model heterogeneity) and train on highly personalized, client-specific data (i.e., data heterogeneity). Although recent work has begun to address client heterogeneity, it mostly considers either model heterogeneity (fang2023robust, ; wu2024fiarse, ; yi2023fedgh, ) or data heterogeneity (chen2024feddat, ; xie2024perada, ; Tamirisa_2024_CVPR, ). Some attempt to tackle both, but often in simplified setups, e.g., assigning each client a model with different ranks of LoRA (hu2022lora, ) adapter while keeping the base architecture identical (cho2024heterogeneous, ; bai2024federated, ), or simulating data heterogeneity by splitting a single dataset into non-i.i.d. partitions and assigning them to different clients (li2019fedmd, ; alam2022fedrolex, ).
To bridge the gap between simulated and real-world federated learning, we propose FedMosaic, a novel multi-modal PFL method that jointly addresses data and model heterogeneity. We reflect real-world data heterogeneity by assigning each client a distinct multi-modal task, e.g., visual question answering and visual reasoning, rather than focusing different clients on non-i.i.d. splits of a single image classification dataset. Moreover, we also account for distribution shifts inherent in the real world, where each client continuously encounters new data over time. To support this, we propose DRAKE, a comprehensive benchmark for multi-modal PFL with 40 diverse tasks. To the best of our knowledge, this is the first benchmark for multi-modal FL that considers data heterogeneity, as well as distribution shifts. We provide an overview of the heterogeneous PFL scenario in Fig. 1.
@@
We then explore effective knowledge sharing across clients that addresses heterogeneous tasks. Previous work shows that naive model averaging (mcmahan2017communication, ) can degrade performance (wu2024personalized, ) due to interference between models trained on heterogeneous tasks (yadav2023ties, ). To reduce interference, inspired by models trained on similar tasks having less conflict (padmakumar2022exploring, ; gurulingan2022curbing, ), we propose SIMilarity-aware Aggregation (SIMA). SIMA constructs a customized global model for each client, based on task similarity. This enables related clients to share knowledge more effectively.
@@
Aggregating model weights enables knowledge sharing without sharing raw data, but only works in the case of identical model architectures between clients (fan2024fedtsa, ). In practice, clients often use different models based on individual hardware constraints or task complexities. Although federated distillation (FD) approaches have been proposed to aggregate heterogeneous models’ knowledge using logits from public data (xie2024perada, ; li2024federated, ), domain discrepancies between public and client data hinder effective knowledge transfer (zhang2023towards, ; wang2023towards, ). Moreover, logit extraction incurs substantial computational costs (malladi2023fine, ), especially for large models. Instead, we introduce PQ-LoRA, which incorporates dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$, whose dimensions depend only on low-rank size $r$, while independent of hidden dimension size, making it shareable among heterogeneous models.
@@
  • Proposing SIMA, a model aggregation strategy that promotes selective knowledge sharing among similar-task models only, addressing data heterogeneity.
@@ 2 Related Work
Federated learning aims to train a strong global model in a distributed manner while preserving privacy by sharing model weights instead of raw data (yurdem2024federated, ; hu2024overview, ). However, global models often fail to converge with heterogeneous client data distributions (i.e., non-i.i.d.) (li2020federated, ; tamirisa2024fedselect, ). To address this challenge, personalized federated learning (PFL) (smith2017federated, ) has emerged alongside the growing importance of model personalization, which allows large foundation models to adapt to individual user preferences, contexts, and needs (shashidhar2024unsupervised, ; zhang2024personalization, ). PFL aims to train a personalized model on each client’s local data by leveraging shared knowledge from other clients to enhance both personalization and generalization (xie2024perada, ).
Data Heterogeneity.
@@
Prior work simulates data heterogeneity by partitioning a single image classification dataset (e.g., MNIST (deng2012mnist, )) into multiple non-i.i.d. subsets, assigning each to a different client. However, assuming such label skew fails to capture real-world data heterogeneity (borazjani2025redefining, ), where clients often tackle different tasks across diverse vision and language domains (madni2024exploiting, ). While  (chen2024feddat, ) makes progress by assigning different VQA datasets to clients, that work remains limited to a single task type (i.e., VQA) and single-image inputs. In contrast, our proposed benchmark, DRAKE, incorporates a broader range of multi-modal tasks, including VQA, visual reasoning, and visual relation detection, covering single- and multi-image inputs, while also modeling temporal distribution shifts within each client, reflecting the evolving and non-stationary nature of real-world data (he2020incremental, ; garg2024tic, ).
Model Heterogeneity.
@@
Federated distillation approaches aggregate knowledge of heterogeneous models using logits from public data (xie2024perada, ; fan2025fedmkt, ). FedMD (li2019fedmd, ) and PerAda (xie2024perada, ) distill averaged logits from local models, while FedMKT (fan2025fedmkt, ) uses those with the lowest loss. However, domain gaps between public and client data hinder knowledge transfer (zhang2023towards, ; wang2023towards, ), while logit sharing raises privacy risks (sun2021vertical, ; lyu2022privacy, ) and incurs significant computational overhead for large models (malladi2023fine, ).
@@
To address the limitation of federated distillation, recent works propose direct model aggregation for heterogeneous models under LoRA-based fine-tuning. HETLoRA (cho2024heterogeneous, ) aligns LoRA modules with varying ranks through zero-padding during aggregation, followed by truncation for redistribution. Similarly, FlexLoRA (bai2024federated, ) multiplies client-specific LoRA matrices to form full matrices, aggregates them, and applies SVD to distribute the aggregated matrix according to each client’s local rank. They address the varying rank sizes, but assume (i) the same input/output dimensions and (ii) identical depths across client models. However, heterogeneous architectures typically have different hidden dimensions and depths (yoshino2023spatially, ; yao2024exploring, ). In contrast, our proposed PQ-LoRA accommodates both different hidden dimensions and varying depths, making it suitable for a more general heterogeneous setup.
@@ 3 Preliminaries
3.1 Low-Rank Adaptation (LoRA)
LoRA (hu2022lora, ) is a widely used parameter-efficient tuning method. It assumes that fine-tuning updates lie in a low-rank space. Building on this assumption, LoRA constrains the weight update $\Delta W$ for a pre-trained weight matrix $W_{p}\in\mathbb{R}^{d_{O}\times d_{I}}$ through low rank decomposition using matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$, where the rank $r\ll\min(d_{O},d_{I})$. During training, only $A$ and $B$ are updated, while $W_{p}$ remains frozen. As a result, when fine-tuning with LoRA, the original output $h_{O}=W_{p}h_{I}\in\mathbb{R}^{d_{O}}$ for an input hidden state $h_{I}\in\mathbb{R}^{d_{I}}$ is modified by incorporating the low-rank update as:
$h_{O}=(W_{p}+\Delta W)h_{I}=(W_{p}+BA)h_{I}.$ (1)
3.2 Problem Statement of Personalized Federated Learning
We consider a PFL setup with $N$ clients, where each client $i\in[N]$ has a local dataset $\mathcal{D}_{i}$. Reflecting real-world scenarios where data arrives incrementally (seo2024learning, ; seo2025budgeted, ), we assume that each client receives a continuous stream of samples $(x^{(i)}_{1},y^{(i)}_{1}),(x^{(i)}_{2},y^{(i)}_{2}),\cdots$. For client models, given a set of model architectures $W=\{W_{1},\dots,W_{K}\}$, client $i$ selects its local pre-trained model $W_{(p,i)}=V(i)$, based on its hardware constraints, where $V$ is a mapping function $V:\{1,\dots,N\}\rightarrow W$. To improve training and communication efficiency, each client trains and shares only its local LoRA adapter $L_{i}$, while freezing the pre-trained model $W_{(p,i)}$, as described in Sec. 3.1. Let $f(W_{(p,i)},L_{i},x)$ denote the forward pass that outputs the logit vector for input $x$, and let the loss function be $\ell(f(W_{(p,i)},L_{i},x),y)$. The empirical loss for client $i$ over its local dataset $\mathcal{D}_{i}$ can be defined as $\mathcal{J}_{i}(L_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(x,y)\in\mathcal{D}_{i}}\ell(f(W_{(p,i)},L_{i},x),y).$ The objective of PFL is then formulated as:
$\min_{\{L_{1},\dots,L_{N}\}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{J}_{i}(L_{i}).$ (2)
@@ 4 Proposed Method
Reflecting the real-world heterogeneity in PFL, we consider both data heterogeneity and model heterogeneity. To address both heterogeneity, we propose FedMosaic, which consists of two key components: SIMA (SIMilarity-aware Aggregation) and PQ-LoRA. SIMA addresses data heterogeneity by reducing interference between local models trained on heterogeneous tasks during aggregation via knowledge sharing among similar-task models only. However, model aggregation (i.e., merging by parameter averaging) becomes infeasible when clients have different architectures. To address model heterogeneity, we introduce PQ-LoRA, which incorporates shareable modules $P\in\mathbb{R}^{r\times r},Q\in\mathbb{R}^{r}$, whose dimensions depend only on low-rank size $r$, while independent of hidden dimension size. We provide an overview of FedMosaic in Fig. 2 and a pseudocode in Sec. A.21.
Figure 2: Overview of our proposed FedMosaic. $L_{i}$ and $G_{i}$ refer to the local and the customized global adapter for the $i_{\text{th}}$ client, respectively. $\hat{g}_{i}$ denotes the decayed last-layer gradient for the $i_{\text{th}}$ client, and $|W|$ the number of layers in model $W$. $\beta_{l}$ is a learnable gating parameter that balances the output from the global and local models in the $l$-th layer. $h$ and $W_{p}$ denote the hidden state input and the pre-trained weight, respectively. $G$ is updated every round and kept frozen during local training.
4.1 Similarity-Aware Aggregation
Model aggregation enables building a single model that excels across multiple tasks without accessing raw data (yadav2024matters, ; wei2025modeling, ). It is thus widely used to construct a shared global model in federated learning. However, aggregating models trained on different tasks often causes parameter interference (yadav2023ties, ; yang2024model, ). Recent work shows models solving similar tasks share more transferable knowledge with fewer conflicts (gurulingan2022curbing, ; padmakumar2022exploring, ). Motivated by this, rather than averaging local models into a single global model, we aggregate them based on task similarity, assigning higher weights to models solving similar tasks and lower weights to less related ones, providing each client with a customized global model.
Client-Wise Gradient $g_{i}$.
To measure task similarity between clients, we calculate the similarity of client-wise gradients. Formally, we calculate the gradient for $i_{\text{th}}$ client as follows:

### v3

# Not All Clients Are Equal: Collaborative Model Personalization on Heterogeneous Multi-Modal Clients

@@ Abstract
As AI becomes more personal, e.g., Agentic AI, there is an increasing need for personalizing models for various use cases. Personalized federated learning (PFL) enables each client to collaboratively leverage other clients’ knowledge for better adaptation to the task of interest, without privacy risks. Despite its potential, existing PFL methods remain confined to rather simplified scenarios where data and models are the same across clients. To move towards realistic scenarios, we propose FedMosaic, a method that jointly addresses data and model heterogeneity with a task-relevance-aware model aggregation strategy to reduce parameter interference, and a dimension-invariant module that enables knowledge sharing across heterogeneous architectures without huge computational cost. To mimic the real-world task diversity, we propose a multi-modal PFL benchmark spanning 40 distinct tasks with distribution shifts over time. The empirical study shows that FedMosaic outperforms the state-of-the-art PFL methods, excelling in both personalization and generalization capabilities under challenging, realistic scenarios.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) with billions of parameters often employ centralized training on massive, heterogeneous datasets using high-performance computing resources (Hurst et al., 2024; Yang et al., 2024a; Team et al., 2024). Such centralized training raises significant concerns about data privacy and high transmission costs (Alvi et al., 2022; Yang et al., 2023). Moreover, as the demand for personalization grows, further fine-tuning is essential to adapt these centrally trained general-purpose models to individual user preferences (Lau et al., 2024; Zhang, 2024).
To address both privacy and personalization, personalized federated learning (PFL) (Smith et al., 2017) has emerged as a decentralized alternative. Recent PFL methods demonstrate that collaboratively leveraging knowledge from other clients significantly improves personalization to the task of interest (Scott et al., 2024; Xie et al., 2024). Unfortunately, despite its potential, most PFL studies overlook the heterogeneity inherent in real-world clients (Zhang et al., 2024b; Bujotzek et al., 2025), where clients typically have distinct models depending on their computational resources (i.e., model heterogeneity) and deal with highly personalized data (i.e., data heterogeneity), as in agentic AI.
@@
Although some recent work addresses client heterogeneity, it mostly considers either model heterogeneity (Fang et al., 2023; Wu et al., 2024a; Yi et al., 2023) or data heterogeneity (Chen et al., 2024; Xie et al., 2024; Tamirisa et al., 2024). Some tackle both, but in less realistic, simplified setups, e.g., assigning each client a model with different LoRA (Hu et al., 2022) adapter ranks while keeping the base architecture identical (Cho et al., 2024; Bai et al., 2024), or lacking data heterogeneity by simply splitting a single dataset into non-i.i.d. partitions (Li & Wang, 2019; Alam et al., 2022).
As a realistic setup tackling both challenges, illustrated in Fig. 1, we consider (i) data heterogeneity, where clients tackle highly personalized tasks, and (ii) model heterogeneity, where clients employ models of different families (e.g., Llama-based (Grattafiori et al., 2024)- vs. Qwen-based (Yang et al., 2024a) MLLMs) and scales (e.g., 1B vs. 3B). To mimic the real-world data heterogeneity, we first introduce DRAKE, a comprehensive benchmark for multi-modal PFL with 40 diverse tasks. Unlike prior works that simulate heterogeneity via non-i.i.d. splits of a single dataset (Xie et al., 2024; Long et al., 2024; Morafah et al., 2024), our benchmark assigns each client a distinct multi-modal task (e.g., visual question answering or visual reasoning), while also incorporating temporal distribution shifts inherent in the real world. To the best of our knowledge, this is the first benchmark for multi-modal FL that considers data heterogeneity, as well as distribution shifts.
Figure 1: Overview of the heterogeneous personalized federated learning scenarios. $L_{i}$ refers to the local model for the $i_{th}$ client. Clients focus on different tasks (i.e., data heterogeneity) where new data are encountered continuously. In addition to data heterogeneity, model architectures may differ across clients (i.e., model heterogeneity) due to differences in hardware constraints.
@@
To address the real-world challenge, we propose FedMosaic that jointly addresses both data and model heterogeneity in the realistic scenario. Under data heterogeneity, naive model averaging (McMahan et al., 2017) often degrades performance (Wu et al., 2024c; Yadav et al., 2023) due to interference between models trained on unrelated tasks. To mitigate the interference, inspired by the fact that models trained on similar tasks have less conflict (Gurulingan et al., 2022), we propose RELevance-guided Aggregation (RELA), which constructs a customized global model for each client based on task-relatedness. This enables related clients to share knowledge more effectively.
@@
Under model heterogeneity, aggregating model weights is infeasible since it requires identical architectures across clients (Fan et al., 2024). Although federated distillation (FD) has been proposed to aggregate heterogeneous models’ knowledge using logits from public data (Xie et al., 2024; Li et al., 2024e), domain discrepancies between public and client data hinder effective knowledge transfer (Wang et al., 2023), and logit extraction is computationally expensive (Malladi et al., 2023), especially for large models. Instead, we propose PQ-LoRA, which incorporates dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$, whose dimensions depend only on low-rank size $r$, while independent of hidden dimension, making them shareable among heterogeneous models.
@@
  • Proposing RELA, a model aggregation strategy that promotes selective knowledge sharing among models learning relevant tasks only, addressing data heterogeneity.
@@ 2 Related Work
Personalized Federated Learning. Federated learning aims to train a strong global model in a distributed manner while preserving privacy by sharing model weights instead of raw data (Yurdem et al., 2024; Hu et al., 2024). With the growing importance of model personalization, which allows large foundation models to adapt to individual user preferences, contexts, and needs (Zhang et al., 2024c), personalized federated learning (PFL) (Smith et al., 2017) has emerged. PFL aims to train a personalized model on each client’s local data by leveraging shared knowledge from other clients to enhance both personalization and generalization while preserving privacy (Xie et al., 2024).
@@
Data Heterogeneity. Prior work simulates data heterogeneity by partitioning a single image classification dataset (e.g., MNIST (Deng, 2012)) into non-i.i.d. subsets per client. However, such label-skew fails to capture real-world data heterogeneity (Borazjani et al., 2025), where clients tackle different tasks across vision and language domains (Madni et al., 2024). Recent work (Chen et al., 2024) moves beyond label skew by assigning different VQA datasets to clients, but remains confined to a single task type and single-image inputs. In contrast, our proposed benchmark, DRAKE, spans a broader range of multi-modal tasks, including VQA, visual reasoning, and visual relation, covering single- and multi-image inputs, while also modeling temporal distribution shifts within each client, reflecting the evolving and non-stationary nature of real-world data (Garg et al., 2024).
@@
Model Heterogeneity. Federated distillation (FD) transfers knowledge across heterogeneous models by sharing logits on public data. FedMD (Li & Wang, 2019) and PerAda (Xie et al., 2024) average logits from local models, while FedMKT (Fan et al., 2025) uses those with the lowest loss. However, public-client domain gaps limit transferability (Wang et al., 2023), and logit sharing introduces both privacy risks (Lyu et al., 2022) and high computational cost for large models (Malladi et al., 2023).
@@
Recent works address the limitations of FD through direct model aggregation for heterogeneous models under LoRA-based fine-tuning. HETLoRA (Cho et al., 2024) and FlexLoRA (Bai et al., 2024) handle varying LoRA ranks through zero-padding/truncation and SVD-based redistribution, respectively. Although they address the varying rank sizes, both assume (i) identical hidden dimensions and (ii) uniform depths across clients, limiting their applicability to heterogeneous architectures that differ in both dimensions and depths (Yao, 2024). In contrast, our proposed PQ-LoRA accommodates both dimensional and depth heterogeneity, enabling more general heterogeneous setups.
@@ 3 Preliminaries
Low-Rank Adaptation (LoRA). LoRA (Hu et al., 2022) assumes that fine-tuning updates lie in a low-rank space. Building on this assumption, LoRA constrains the weight update $\Delta W$ for a pre-trained weight matrix $W_{p}\in\mathbb{R}^{d_{O}\times d_{I}}$ through low rank decomposition using matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$, where the rank $r\ll\min(d_{O},d_{I})$. During training, only $A$ and $B$ are updated, while $W_{p}$ remains frozen. With LoRA, the original output $h_{O}=W_{p}h_{I}\in\mathbb{R}^{d_{O}}$ for an input hidden state $h_{I}\in\mathbb{R}^{d_{I}}$ is modified by incorporating the low-rank update as $h_{O}=(W_{p}+\Delta W)h_{I}=(W_{p}+BA)h_{I}$.
Problem Statement of Personalized Federated Learning. We consider a PFL setup with $N$ clients, where each client $i\in[N]$ has a local dataset $\mathcal{D}_{i}$. Reflecting real-world scenarios where data arrives incrementally (Seo et al., 2024; 2025), we assume that each client receives a continuous stream of samples $(x^{(i)}_{1},y^{(i)}_{1}),(x^{(i)}_{2},y^{(i)}_{2})$$\cdots$. Given a set of model architectures $W=\{W_{1},\dots,W_{K}\}$, each client $i$ selects its local pre-trained model $W_{(p,i)}=V(i)$, based on its hardware constraints, where $V$ is a mapping function $V:\{1,\dots,N\}\rightarrow W$. For efficiency, each client trains and shares only its local LoRA adapter $L_{i}$, while freezing $W_{(p,i)}$. Let $f(W_{(p,i)},L_{i},x)$ denote the forward pass that outputs logits for input $x$, and let the loss function be $\ell(f(W_{(p,i)},L_{i},x),y)$. The empirical loss for client $i$ over its local dataset $\mathcal{D}_{i}$ can be defined as $\mathcal{J}_{i}(L_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(x,y)\in\mathcal{D}_{i}}\ell(f(W_{(p,i)},L_{i},x),y).$ The objective of PFL is then formulated as $\min_{\{L_{1},\dots,L_{N}\}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{J}_{i}(L_{i})$.
@@ 4 Proposed Method
To address the real-world heterogeneities, i.e., data heterogeneity and model heterogeneity, that hinder client collaborations in personalized federated learning, we propose FedMosaic, illustrated in Fig. 2, comprising of: RELA (RELevance-guided Aggregation) and PQ-LoRA. RELA mitigates data heterogeneity by restricting knowledge sharing to local models trained on related tasks, thus reducing interference during aggregation (i.e., merging by parameter averaging). Under model heterogeneity, model aggregation becomes infeasible. To aggregate different models, we introduce PQ-LoRA, which incorporates shareable modules $P\in\mathbb{R}^{r\times r},Q\in\mathbb{R}^{r}$ in LoRA, whose dimensions depend only on low-rank size $r$, not on the hidden dimension size. We provide a pseudocode in Sec. A.28.
4.1 Relevance-Guided Aggregation
Model aggregation builds a single model that excels in multiple tasks without accessing raw data (Wei et al., 2025). It is thus widely used to construct a shared global model in federated learning. However, naively averaging models trained on different tasks often causes parameter interference (Yadav et al., 2023; Yang et al., 2024b). Recent work shows that models solving similar tasks share more transferable knowledge with fewer conflicts (Gurulingan et al., 2022). Motivated by this, we replace the uniform averaging with a relevance-guided strategy that assigns a higher aggregation weight to clients with closer task relations, providing each client with a customized global model.
Figure 2: Overview of proposed FedMosaic. On every round, the local PQ-LoRA $L_{i}$ fine-tuned during local training and the sanitized last layer gradient $\tilde{g_{i}}$ are uploaded from the $i_{\text{th}}$ client to server. The last layer gradient $g$ is extracted from the small pre-trained model $W_{s}$, which is then EMA updated to $\hat{g}$ and then compressed to $\tilde{g}$, sequentially. Note that the gradient computation is performed every $m$ iterations. In server, the sanitized gradients $\tilde{g_{i}}$ are used to measure client task relevance and to build customized global PQ-LoRA $G_{i}$, which is distributed and kept frozen. $h$ and $W_{p}$ denote the hidden state input and the pre-trained weight, respectively. $\beta$ is a learnable gating parameter that balances the output from the global and local models, and $|W_{p}|$ is the number of layers in the model.
Client-Wise Gradient $g_{i}$. To measure task relevance between clients, we calculate the similarity of client-wise gradients. Formally, we calculate $g_{i}$, the gradient for $i_{\text{th}}$ client as follows: